A 20g sample of P4O10 contains how many molecules of phosphorus pentaoxide?
1
Expert's answer
2019-05-15T05:47:47-0400

1. Find moles of "P_4O_{10}"


"n=\frac{m}{M}=\frac{20g}{283.89\frac{g}{mol}}=0.070 mol"


As 1 molecule of "P_4O_{10}" cobtains 2 molecules of "P_2O_5" , then moles of "P_2O_5" :

"n(P_2O_5)=n(P_4O_{10})\times 2 = 0.070\times 2= 0.14 mol"

2.Find number of molecules of "P_2O_5"


"N=N_A\times n=6.022\times10^{23}\times 0.14 =8.4\times 10^{22} molecules"